FEL Upgrade Project Weekly Brief
March 19-23, 2001


This week the FEL project had the pleasure of hosting an important visit on Thursday:  Vice Admiral G. P. Nanos, Rear Admiral M. Mathis from the Naval Sea Systems Command, and Brig. General  R. Natonski (USMC).  The  group was impressed with the FEL project to date and was particularly interested in the capabilities of the existing FEL user facility for laser-materials and propagation studies of interest to the Navy.  We were told to expect a return visit and we look forward to that opportunity.


FEL management attention this week was focussed on preparing for the visit of the above noted dignitaries on Thursday and reviewing our preparations for next week's semiannual review of the FEL upgrade project.  The project has progressed well since the last review despite the stress on all of the project staff with concurrent FEL operations.  We look forward to presenting our status report to the review committee next week.

WBS 3 (Beam Physics):

Paperwork was generated for the Rev 1.1 ECR; review preparations were undertaken. Discussions were held about diagnostic systems layout. Archival DIMAD design files for the IR (032101/032101out) and UV
(uv0313/uv313out) Revision 1.1 driver design were generated in preparation for design documentation.
Injection line dipole specifications were analyzed and will be documented in a forthcoming technical note.

WBS 4 (Injector):

Continued work on the assembly drawings for the shield and internal cesiation design, these will be complete early next week.

Gun HVPS - The spare unit is at the vendor for the upgrade.  We are designing the extension to the HVPS Tank and the mounting of the 2nd multiplier stack in the tank.  This effort has just started.

WBS 5 (SRF):

Received 5 helium vessels from PHPK.
Post end can award visit to Ability Engineering completed.
Completed welding of one 7 cell cavity.

WBS 6 (RF):

Zone 3 - No progress.  Waiting on parts.

Zone 4 - Down to last few items in the punch list.  Circulator for position 8 being repaired and is due by 7/1/01.  RF Control Module in position 8 needs calibration coefficients reloaded.  The Mod Anode Monitoring and the Cavity  Waveguide Window Temperature (CWWT) systems need to be checked.  The checking and updating the documentation for the Cross Connect wiring needs to be completed.  It is about 50 % complete.

Injector RF - The first article 100 kW Klystron is now scheduled for factory acceptance testing on 6/20/01.  The upgrade of the HVPS is still being reviewed by the vendor.  I am supplying him different design options that would be acceptable to JLab.  We may have to do this work in-house.

WBS 8 (I&C):

Sixteen (16) BPMs (of 44) have been TDR'd and are in final welding. The welded bellows for the 36 3" shielded beam viewers are also in, these were late by 3 weeks from Standard Bellows.  These are being welded on to the viewers. Both of these fab jobs are being done by Master Machine in NN. They are doing very well and are holding to the schedule.

James and Dan are installing the new LSS box in Lab 2. This will allow for operation of the mirror test stand. Once lab 2 is done they will move on to lab 5 than all labs will be fully commissioned.

Vacuum design and procurements continue for the upgrade. We have gathered, or ordered, most of the parts for the new Ion Pump Power Supplies. Requests for quotes for the fabrication of the rest of the IPPS have been sent out.  No figures yet.  The drawing for the faceplate has been revised to relax the tolerances on the holes.  Drawings for the User power supply boxes are complete.  We are still waiting for some parts before we can get the new ones made.

An IOC is now working in lab 5 as a VME test stand. This will be used to test and commission all VME hardware prior to installation in the accelerator. It now has a fully tested GPIO card. A computer (cast-off from AES) was reworked and commissioned as a dedicated Altera programming station, this is available for general (labwide) use and is located in the back of room 216

There is continued progress on the GPIO card schematics (~90%) as well as other drawing for both "as-builts" and for upgrade efforts. We received the face-plates that were sent to Service Metals for the GPIB and GPVS systems. Mike is adding pictures of the zones and racks to the web site. We have begun work on a new CD that will be used as an intro to the capabilities of the FEL, this will include specs and building/lab layout as well as the much enjoyed AVIs! Everyone's assistance is needed for this to be representative of our successes.  Drawing numbers will follow as they are completed.  Final preparations for the web server 'laser' upgrade.

WBS 9 (Transport):

Injector Dipoles (DU/DV)
o  DULY Research continues to do small manipulations to their 3D model (in RADIA) of the GV dipole but
    with the increased number of nodes.  They have achieved the flatness specification and are working on the
    correctly uniform field integral change.  Their drawing set and the drafting model of the GV (Small injector
    dipole) is being worked on.
Optical Chicane Dipole (DW)
o  We continued back check of the remaining parts, the package should be ready in one week.
Arc Dipoles (GY, GX, GQ)
o  AES continues progress in designing 180 degree dipole (GY) as they are cleaning up numerous yoke details
    and are laying out the leads for the path length correctors.
o  AES, confidant in the resolution of their model, is refining. They have had to increase the width of the path
    length corrector coils to maintain field flatness through the ends and make the slot in the pole face (for the
    path length correctors 0 wedge shaped to maintain a field integral that increases correctly with radius.
QX (Trim Quad)
o We started further analysis of the magnetic model
o We started laying out the girders of the Return Leg.  These positions are the most general of the quad
    girders. We will learn some design techniques from this exercise that allow us to design the more complex
    regions more rapidly.
o We had feed back on the steel specification from Bethlehem Lukens, indicating they could try to obtain our
    tight metallurgical specifications but could only guarantee slightly higher values.  They also will not be able to
    make the slabs in one heat.  We modified the specification and it is now being checked by Leigh Harwood.

WBS 10 (Wiggler):

A final check of the remaining drawings for the optical klystron vacuum chamber and diagnostics was completed.  Drawings should be ready for sign-off shortly.  A task order to find the optimum end corrector strength for the wigglers was drafted and sent to AES.  A clogged cooling line in the dispersion section was found and  fixed.  The poles are slightly askew so the jaws are being shimmed to take out the angle. We are on track to begin measurements of the magnet next week.

We fixed a plugged braise joint in the Dispersion Section for the Optical Klystron that was missed by the coil vendor's inspection.  The gap is up to .002 inch too wide at the top.  We will shim the mounting yoke to close the gap and set the magnet on the test stand next week.

WBS 11 (Optics):

The bid package for two of our prototype deformable mirrors is in Procurement.  Preparations to ready User Lab 2 for cavity mirror tests is proceeding at a rapid pace; the Nd:YAG laser that will used for initial tests (a gift from AK Steel) is installed, and test run (set up to not lase).  We will do a test pump down of the vacuum vessel today (3/23).  The I. & C. group is installing the LSS hardware to support use of either this laser or the FEL in the lab.  We have new drawings and placement of components of the Upgrade optical cavity, they will be circulated for comment.  New nomenclature for some of these components was created.  Metal shields, used to protect the  windows that send beam into the user labs were received from our shop, as were some parts for the manual beam shutters.

We had a service visit for the drive laser.  The vendor (Coherent) will declare the laser obsolete at the end of this month, so this was our last opportunity to have this service performed.  We now have improved performance (4W of SHG) from our well-used Nd:YLF laser rods, and a spare, new rod that produces > 7W of SHG.  We purchased spare components that should keep the laser operational for many more years.  A valve assembly for the drive laser heat exchanger began leaking building HVAC water, and will be replaced next week.


No operations this week.  We continued preparations in Lab 2 for the start-up of the mirror test stand for the
10 kW Upgrade.

Privacy and Security Notice